OpenStudy (anonymous):

Choose the correct simplification of 9x^2(4x + 2x^2 – 1) PLEASE EXPLAIN !!!!!!

OpenStudy (anonymous):

multiply each term in the second parentheses by \(9x^2\)

OpenStudy (anonymous):

\[9x^2\times 4x+9x^2\times 2x^2-9x^2\] is a start

OpenStudy (anonymous):

thats all ?

OpenStudy (anonymous):

no, now you have to compute what you get

OpenStudy (anonymous):

for example \(9\times 4=36\) and \(x\times x^2=x^3\) so teh first term is \[36x^3\]

OpenStudy (lgbasallote):

hmm...wouldn't simplifying be factoring?

OpenStudy (anonymous):

okay and then you put all the terms in one equation right

OpenStudy (anonymous):

@satellite73 36x4 + 18x3 − 9x2 answer??